# Constraining anti-baryonic dark matter through correlated nucleon decay signatures

### v1

@@ V Summary
Baryon number violation is a key ingredient in generating the observed baryon asymmetry in the universe. Since the Standard Model conserves baryon number at the perturbative level, observing these rare decay processes would provide clear evidence of New physics. On the other hand, current experimental constraints place the proton/neutron lifetime above $10^{34}\hskip 2.84544pt\text{years}$, setting stringent limits on models with baryon-violating processes. Complementary to this, di-nucleon decay, which involves the simultaneous decay of two bound nucleons, provides an alternative probe of baryon number violation. Processes that produce back-to-back Cherenkov radiations, such as $pp\to\pi^{+}\pi^{+}$, $nn\to\pi^{0}\pi^{0}$ etc, are actively searched by current experiments. The operators that generate such di-nucleon processes are interesting as their Wilson coefficients are less suppressed. In addition to visible decay channels, invisible neutron decay like $n\to\bar{\nu}+\text{invisible}$, and $nn\to\bar{\nu}\bar{\nu}$, provide another potential signature of baryon number violation. Experimental searches for unstable nuclei in large-volume detectors aim to detect such decays, which, if observed, could provide crucial insights into the nature of the dark sector and baryon number-violating interactions. Models that predict such dark matter induced baryon number violating interactions Davoudiasl:2010am ; Davoudiasl:2011fj ; Huang:2013xfa are of interest because they bring the scale of effective field theory within the collider limits.
@@
In this article, we explore how di-nucleon decay is generated when spontaneous breaking of baryon number is included. While the spontaneous breaking of baryon number in the dark sector has been of interest in the context of dark CP violation and dark phase transitions, we observe in our paper that induced nucleon decay generated by anti-baryonic dark matter and di-nucleon decays are correlated. Our analysis demonstrates that the effective field theory (EFT) operators responsible for nucleon decay ($n\to\bar{\nu}+\text{invisible},\ n\to\pi^{0}+\text{invisible}$) also generate di-nucleon decay processes ($nn\to\bar{\nu}\bar{\nu},\ nn\to\pi^{0}\pi^{0}$) at the one-loop level. This connection provides a complementary approach to probe induced baryon number violation, which plays a crucial role in understanding the origin of baryon asymmetry and its possible link to dark matter.
@@
Based on current experimental constraints from KamLand and Super-Kamiokande for di-nucleon decay, the analysis constrains the new physics scale of the nucleon decay operator in the range of $10^{6}-10^{8}\hskip 2.84544pt\mathsf{GeV}$, depending on the specific decay mode. These limits are competitive with those derived from traditional nucleon decay searches, reinforcing the viability of di-nucleon decay as a complementary and potentially more accessible probe of baryon number violation, especially since it predicts a back-to-back Cherenkov radiation signature.
@@
For the $B-L$ preserving decay channel, the fraction of anti-baryonic dark matter density in the total dark matter flux at Earth is computed to be $\rho_{\text{DM}}=0.99\times 10^{-4}\hskip 2.84544pt\mathsf{GeV}/\mathsf{cm}^{3}$ for dark matter with mass $M_{\text{DM}}=10\hskip 2.84544pt\mathsf{GeV}$. In contrast, for the $B-L$ violating decay, the dark matter density becomes $\rho_{\text{DM}}=0.22\times 10^{-3}\hskip 2.84544pt\mathsf{GeV}/\mathsf{cm}^{3}$. A higher fraction of the total flux at Earth is allowed, provided the new physics scales are very high. These values indicate that neutron decay channels provide an indirect probe of the local dark matter environment and can impose limits on the anti-baryonic dark matter density. On the other hand, if we assume that this dark matter accounts for the entire flux on Earth, the upper limit on the mass of dark matter is strongly constrained to be approximately $55\hskip 2.84544pt\mathsf{GeV}$ and $80\hskip 2.84544pt\mathsf{GeV}$ for the $B-L$ preserving and $B-L$ violating scenarios, respectively.
@@ Acknowledgments
M.T.A. acknowledges the financial support of DST through the INSPIRE Faculty grant DST/INSPIRE/04/2019/002507.

### v2

@@ V Summary
Baryon number violation is a key ingredient in generating the observed baryon asymmetry in the universe. Since the Standard Model conserves baryon number at the perturbative level, observing these rare decay processes would provide clear evidence of New Physics. On the other hand, current experimental constraints place the proton/neutron lifetime above $10^{34}\hskip 2.84544pt\text{years}$, setting stringent limits on models with baryon-violating processes. Complementary to this, di-nucleon decay, which involves the simultaneous decay of two bound nucleons, provides an alternative probe of baryon number violation. Processes that produce back-to-back Cherenkov radiations, such as $pp\to\pi^{+}\pi^{+}$, $nn\to\pi^{0}\pi^{0}$ etc, are, hence, actively searched by current experiments. The operators that generate such di-nucleon processes are interesting as their Wilson coefficients are less suppressed. In addition to visible decay channels, invisible neutron decay like $n\to\bar{\nu}+\text{invisible}$, and $nn\to\bar{\nu}\bar{\nu}$, provide another potential signature of baryon number violation. Experimental searches for unstable nuclei in large volume detectors aim to detect such decays, which, if observed, could provide crucial insights into the nature baryon number violation.
@@
In this article, we explore how di-nucleon decay is generated with spontaneous breaking of baryon number in the dark sector. While the spontaneous breaking of baryon number in the dark sector has been of interest in the context of dark CP violation and dark phase transitions, we observe in our paper that induced nucleon decay generated by anti-baryonic dark matter and di-nucleon decays are correlated. Our analysis demonstrates that the effective field theory (EFT) operators responsible for induced nucleon decay ($n\to\bar{\nu}+\text{invisible},\ n\to\pi^{0}+\text{invisible}$) also generate di-nucleon decay processes ($nn\to\bar{\nu}\bar{\nu},\ nn\to\pi^{0}\pi^{0}$) at the one-loop level. This connection provides a complementary approach to probe induced baryon number violation, which plays a crucial role in understanding the origin of baryon asymmetry and its possible link to dark matter.
@@
Based on current experimental constraints from KamLand and Super-Kamiokande for di-nucleon decay, the analysis constrains the new physics scale of the induced nucleon decay operator in the range of $\mathcal{O}(10^{7}\mathsf{GeV})$, depending on the specific decay mode. These limits are competitive with those derived from traditional nucleon decay searches, reinforcing the viability of di-nucleon decay as a complementary and potentially more accessible probe of baryon number violation, especially since it predicts a back-to-back Cherenkov radiation signature.
@@
For the BLB-L preserving decay channel, the fraction of anti-baryonic dark matter density in the total dark matter flux at Earth is computed to be $\rho_{\text{DM}}=0.2\ \mathsf{GeV/cm^{3}}$ for dark matter with mass $M_{\text{DM}}=2\mathrm{GeV}$ and $R_{\bar{\nu}\bar{\nu}}=10^{-5}$. In contrast, for the BLB-L violating decay, the dark matter density becomes $\rho_{\text{DM}}=0.3\times 10^{-2}\ \mathsf{GeV}/\mathsf{cm}^{3}$ for a similar value of $R_{\pi\pi}$. These values indicate that neutron decay channels provide an indirect probe of the local dark matter environment and can impose limits on the anti-baryonic dark matter density. On the other hand, if we assume that this dark matter accounts for the entire flux on Earth, the upper limit on the mass of dark matter is strongly constrained to be approximately $14\ \mathsf{GeV}$ and $20\ \mathsf{GeV}$ for the BLB-L preserving and BLB-L violating scenarios, respectively.
@@ Acknowledgments
M.T.A. acknowledges the financial support of DST through the INSPIRE Faculty grant DST/INSPIRE/04/2019/002507. A.B.K. thanks IISER Thiruvananthapuram for hosting her during this project.